You have a relative who has accumulated savings of kshs.250, 000 over his working lifetime and now plans to retire. Assuming that he wishes to withdraw equal installments from these savings for the next 25 years of his life, how much will each installment amount to if he is earning 5% on his savings
1
Expert's answer
2019-04-03T09:03:32-0400

Using the annuity payment formula:

"P = r*PV\/(1 - (1 + r)^{-n}),"

"P = 0.05*250,000\/(1 - (1 + 0.05)^{-25}) = 17,738.11."